# Testing Refactoring Engine via Historical Bug Report driven LLM

### v1

@@ I Introduction
Despite the active development of refactoring engines, they can be buggy. These bugs can silently change the program behaviors, produce uncompilable programs, or induce inconsistencies in the modified code [10]. To ensure the reliability of refactoring engines, several techniques have been proposed to identify Refactoring Engine Bugs (REBs) via input programs generation, including (1) template-based input program generation technique [11] that relies on manually crafted imperative generators, and (2) SAFEREFACTOR-based tools [12, 13, 14, 15, 16, 17, 18] that relies that random test generation. However, there are several limitations in these tools that hinder their effectiveness in identifying REBs. Firstly, template-based input program generation techniques like ASTGen [11] usually require developers to manually write a set of predefined templates for generating input programs. However, developers may not have sufficient knowledge nor insights about the types of program structures that are more likely to trigger bugs. Moreover, modern refactoring engines in IDE like IntelliJ IDEA support a variety of refactoring types (e.g., Type Migration, Rename, Extract Variable, Extract Method, Extract Parameter, and Change Method Signature) which makes manually designing templates labor-intensive and impractical. Secondly, SAFEREFACTOR-based tools [12, 13, 14, 15, 16, 17] rely on automatic test generation tools to identify refactoring bugs but these approaches may be ineffective because: (1) they rely mosly on random input generation which did not consider the bug-triggering conditions for the input programs (e.g., the input programs for the Rename Class refactoring need to contain an inner class with a specific name), thus leading to low effectiveness, and (2) they only generate test cases for the unchanged common methods before and after refactoring in the project, neglecting the impact of refactoring on the deleted or newly added methods after the refactoring. Indeed, a study [19] revealed that more than 50% of refactoring engine bugs fail to be detected by existing automatic test generation tools (i.e., Randoop [20] used in SAFEREFACTOR, and EvoSuite [21]).
@@ III-E Mutation Strategy
Figure 3 bottom shows one variant generated by the LLM using the prompt template in Table V whose bug-triggering characteristic is set to Java anonymous class. This variant is refactoring-preserving, which means the “Pull Up” refactoring could still be applied on the “methodToBePulledUp()” method on variant. This variant successfully triggered one bug in the newest version of Eclipse (2024-09), resulting in an uncompilable refactored program. We have reported the bug together with the variant input program and bug reproduce steps to the Eclipse’s issue tracker systems [32].
@@ III-F Differential Testing
We manually refactor the variants in the newest version of refactoring engines by applying the same refactoring as in the seed bug reports, and compare the refactoring results of Eclipse and IntelliJ IDEA using differential testing. Specifically, based on prior work on testing refactoring engines [11], we adopt the following oracles:
@@ IV-A Implementation
As shown in Table VI, we randomly select five historical bug reports from our seed dataset for our experiment. Specifically, two seeds are from Eclipse and three are obtained from IntelliJ IDEA. These seeds cover five refactoring types and two symptoms. We use OpenAI API [47] to programmatically invoke the ChatGPT. The model type is gpt-4o-mini with the default settings [48] (e.g., temperature is 1) and the knowledge cut-off date is Oct 2023. For each seed, we iterate the characteristics listed in Table IV, and for each characteristic, we ask the LLM to generate ten variants. In conclusion, we generate 30 variants for each seed considering three different characteristics. In total, 150 variants are generated for the five seeds in Table VI. We set the JDK version in Table V to 22.0.1, and the corresponding refactoring definition is obtained from [49]. We manually tested the input program variants generated by our tool in the latest version of Eclipse (2024-09) and IntelliJ IDEA (2024.2.4). All experiments are run on a workstation with 2.6GHz 6-core Intel Core i7 CPU and Windows 10, 64-bit operating system.
@@ V-A RQ1: Effectiveness
ID IDE Issue No. Refactoring Type Symptom Status B-1 Eclipse 1785 Extract Method Compile error Submitted B-2 Eclipse 1824 Make Static Compile error Confirmed B-3 Eclipse 1783 Inline Method Compile error Submitted B-4 Eclipse 1781 Inline Method Compile error Submitted B-5 Eclipse 1780 Inline Method Compile error Fixed B-6 Eclipse 1779 Inline Method Compile error Submitted B-7 Eclipse 1778 Inline Method Compile error Submitted B-8 Eclipse 1777 Pull Up Compile error Submitted B-9 Eclipse 1776 Pull Up Compile error Submitted B-10 Eclipse 1775 Pull Up Compile error Submitted B-11 Eclipse 1774 Pull Up Failed refactoring Submitted B-12 Eclipse 1773 Pull Up Compile error Fixed B-13 Eclipse 1772 Pull Up Compile error Submitted B-14 Eclipse 1766 Pull Up Compile error Submitted B-15 Eclipse 1823 Pull Up Compile error Confirmed B-16 IDEA 364110 Pull Members Up Compile error Confirmed B-17 IDEA 362805 Pull Members Up Compile error Confirmed B-18 IDEA 362804 Pull Members Up Compile error Confirmed
@@ IX Conclusion
In this paper, we introduced a novel approach to test refactoring engine based on historical bug reports and bug-triggering input program characteristics empowered by Large Language Model (LLM). RETester does not require manually designing templates and can easily generalize to diverse refactoring types. The experimental results show that RETester can detect various new bugs in both Eclipse and IntelliJ IDEA. In total, we identified 18 new bugs in the latest version of both refactoring engines. By the submission time of our paper, seven bugs were confirmed, three bugs have already been fixed.

### v2

@@ I Introduction
Despite the active development of refactoring engines, they can be buggy. These bugs can silently change the program behaviors, produce uncompilable programs, or induce inconsistencies in the modified code [10]. To ensure the reliability of refactoring engines, several techniques have been proposed to identify Refactoring Engine Bugs (REBs) via input programs generation, including (1) template-based input program generation technique [11] that relies on manually crafted imperative generators, and (2) SAFEREFACTOR-based tools [12, 13, 14, 15, 16, 17, 18] that relies that random test generation. However, there are several limitations in these tools that hinder their effectiveness in identifying REBs. Firstly, template-based input program generation techniques like ASTGen [11] usually require developers to manually write a set of predefined templates for generating input programs. However, developers may not have sufficient knowledge nor insights about the types of program structures that are more likely to trigger bugs. Moreover, modern refactoring engines in IDE like IntelliJ IDEA support a variety of refactoring types which makes manually designing templates labor-intensive and impractical. Secondly, SAFEREFACTOR-based tools [12, 13, 14, 15, 16, 17] rely on automatic test generation tools to identify refactoring bugs without considering the bug-triggering ability of the input programs (e.g., input programs contain lambda expression or anonymous class are more error-prone during refactoring), thus leading to low effectiveness.
@@ III-E Mutation Strategy
Figure 3 bottom shows one variant generated by the LLM using the prompt template in Table V whose bug-triggering characteristic is set to Java anonymous class. This variant is refactoring-preserving, which means the “Pull Up” refactoring could still be applied on the “methodToBePulledUp()” method on variant. This variant successfully triggered one bug in the newest version of Eclipse (2024-09), resulting in an uncompilable refactored program. We have reported the bug together with the variant input program and bug reproduce steps to the Eclipse’s issue tracker systems [29]. We measure the LLM’s ability to perform RPT in RQ1.
@@ III-F Differential Testing
We manually refactor the variants in the newest version of refactoring engines by applying the same refactoring as in the seed bug reports, and compare the refactoring results of Eclipse and IntelliJ IDEA using differential testing. To ensure the reliability and objectivity of the manual process, two annotators independently refactor each variant in the IDEs according to the reproduce procedures in the seed bug reports. Then, two annotators hold a meeting to resolve their disagreements. Based on prior work on testing refactoring engines [11], we adopt the following oracles:
@@ IV-A Implementation
As shown in Table VI, we randomly select five historical bug reports from our seed dataset for our experiment. Specifically, two seeds are from Eclipse and three are obtained from IntelliJ IDEA. These seeds cover five refactoring types and two symptoms. We use OpenAI API [44] to programmatically invoke the ChatGPT. The model type is gpt-4o-mini with the default settings [45] and the knowledge cut-off date is Oct 2023. For each seed, we iterate the characteristics listed in Table IV, and for each characteristic, we ask the LLM to generate ten variants. In conclusion, we generate 30 variants for each seed considering three different characteristics. In total, 150 variants are generated for the five seeds in Table VI. We set the JDK version in Table V to 22.0.1, and the corresponding refactoring definition is obtained from [46]. We manually tested the input program variants generated by our tool in the latest version of Eclipse (2024-09) and IntelliJ IDEA (2024.2.4). All experiments are run on a workstation with 2.6GHz 6-core Intel Core i7 CPU and Windows 10, 64-bit operating system.
@@ V-A RQ1: Effectiveness
ID IDE Issue No. Refactoring Type Symptom Status B-1 Eclipse 1785 Extract Method Compile error Submitted B-2 Eclipse 1824 Make Static Compile error Confirmed B-3 Eclipse 1783 Inline Method Compile error Submitted B-4 Eclipse 1781 Inline Method Compile error Submitted B-5 Eclipse 1780 Inline Method Compile error Fixed B-6 Eclipse 1779 Inline Method Compile error Submitted B-7 Eclipse 1778 Inline Method Compile error Submitted B-8 Eclipse 1777 Pull Up Compile error Submitted B-9 Eclipse 1776 Pull Up Compile error Submitted B-10 Eclipse 1775 Pull Up Compile error Submitted B-11 Eclipse 1774 Pull Up Failed refactoring Submitted B-12 Eclipse 1773 Pull Up Compile error Fixed B-13 Eclipse 1772 Pull Up Compile error Submitted B-14 Eclipse 1766 Pull Up Compile error Submitted B-15 Eclipse 1823 Pull Up Compile error Fixed B-16 IDEA 364110 Pull Members Up Compile error Confirmed B-17 IDEA 362805 Pull Members Up Compile error Confirmed B-18 IDEA 362804 Pull Members Up Compile error Confirmed
@@ IX Conclusion
In this paper, we introduced a novel approach to test refactoring engine based on historical bug reports and bug-triggering input program characteristics empowered by Large Language Model (LLM). RETester does not require manually designing templates and can easily generalize to diverse refactoring types. The experimental results show that RETester can detect various new bugs in both Eclipse and IntelliJ IDEA. In total, we identified 18 new bugs in the latest version of both refactoring engines. By the submission time of our paper, seven bugs were confirmed, three of them were fixed.
@@
Acknowledgments
This work is supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) Discovery Grants RGPIN-2024-04301.